The probability that at least one of A and B occurs is 0.6 . If A and B occur simultaneously with probability 0.3, then P(A ) P(B ) ' ' +
is
(A) 0.9 (B) 0.15 (C) 1.1 (D) 1.2

Grade:

2 Answers

Arun
3 years ago
P (AUB)=0.6,P (A intersection B)=0.2
P (AUB)=P (A)+P (B)-P (A intersection B)
0.6=P (A)+P (B)-0.2
0.6+0.2=P (A)+P (B)
P (A)+P (B)=0.8
P (A bar )+P (Bbar)=(1-p (A))+(1-p (B))
=2-(P(A)+P (B))
=2-(0.8)
=1.2
 
Vikas TU
3 years ago
P (AUB)=0.6,P (A intersection B)=0.2
P (AUB)=P (A)+P (B)-P (A intersection B)
0.6=P (A)+P (B)-0.2
0.6+0.2=P (A)+P (B)
P (A)+P (B)=0.8
P (A bar )+P (Bbar)=(1-p (A))+(1-p (B))
=2-(P(A)+P (B))
=2-(0.8)